Edna Recording Studios, Inc., reported earnings available to common stock of $4200000 last year. From these earnings, the company paid a dividend of $1.26 on each of its 1000000 common shares outstanding. The capital structure of the company includes 40% debt, 10% preferred stock, and 50% common stock. It is taxed at a rate of 40%. The company currently has $1000-par-value, 10% coupon, 5-year bonds that can be sold for $1175 each. What is the after tax cost of debt financing? (Show your work/calculations)

A. 5.86%

B. 6.99%

C. 3.52%

D. 10.00%
Answer #1

We can calculate the before tax cost of debt (I/Y) using a financial calculator

Following data should be used as inputs.

FV = $1,000, PMT = 10% * $1,000 = $100, N = 5, PV = $ -1175

Using 1175 as a negative amount because that's what an investor would pay for and thus indicates an outflow.

Using these inputs in the financial calculator, we get I/Y = 5.86%

Now the after tax cost of debt = I/Y * (1 - Tax) = 5.86% * (1-40%) = 5.86% * 0.6 = 3.52%

Option c is correct